An employee who produces x units per hour earns an hourly wage of y = 0.55x + 7 (in dollars). Find the hourly wage for an employee who produces 13 units per hour.
Thank you

Answer by stanbon(75887)

An employee who produces x units per hour earns an hourly wage of y = 0.55x + 7 (in dollars). Find the hourly wage for an employee who produces 13 units per hour.
-----------
Substitute 13 for "x" to solver for y=f(13)
f(13)=0.55*13 + 7
f(13)=7.15+15
f(13)=$22.15
-----------
Cheers,
Stan H.